# Learning to Receive Help: Intervention-Aware Concept Embedding Models

### v1

@@ 1 Introduction
It is important to know how to ask for help, but also important to know how to receive help. Knowing how to react to feedback positively allows for bias correction Gusukuma et al. (2018) and efficient learning Webb (1995) while being instrumental for mass collaborations Doan et al. (2010). Nevertheless, although the uptake of feedback is ubiquitous in real-world decision-making, the same cannot be said about modern artificial intelligence (AI) systems, where deployment tends to be in isolation from experts whose feedback could be queried.
Progress in this aspect has recently come from Explainable AI, where interpretable Deep Neural Networks (DNNs) that can benefit from expert feedback at test-time have been proposed Koh et al. (2020); Espinosa Zarlenga et al. (2022); Havasi et al. (2022); Magister et al. (2022). In particular, Concept Bottleneck Models (CBMs) Koh et al. (2020), a family of interpretable DNNs, enable expert-model interactions by generating, as part of their inference process, explanations for their predictions using high-level “concepts”. Such explanations allow human experts to better understand a CBM’s prediction based on interpretable units of information (e.g., input is a “cat” because it has “paws” and “whiskers”) rather than low-level features (e.g., input pixels). This design enables humans to provide feedback to the CBM at test-time via concept interventions Koh et al. (2020) (Figure 1), a process specific to CBM-like architectures in which an expert analyses a CBM’s inferred concepts and corrects mispredictions. Concept interventions allow a CBM to update its final prediction after incorporating new expert feedback, and have been shown to lead to significant performance boosts Koh et al. (2020); Espinosa Zarlenga et al. (2022); Shin et al. (2023); Chauhan et al. (2022). Nevertheless, and in contrast to these results, recent works have found that concept interventions can potentially increase a CBM’s test-time error depending on its choice of architecture Espinosa Zarlenga et al. (2023) and its set of train-time annotations Espinosa Zarlenga et al. (2022). This discrepancy suggests that a CBM’s receptiveness to interventions is neither guaranteed nor fully understood.
@@
In this paper, we posit that the sensitivity of a CBM’s receptiveness to interventions is an artefact of its training objective function. Specifically, we argue that CBMs lack an explicit incentive to be receptive to interventions as they are neither exposed to interventions at train-time nor optimised to perform well under interventions. To address this, we propose a novel intervention-aware objective function and architecture based on Concept Embedding Models (CEMs) Espinosa Zarlenga et al. (2022), a generalisation of CBMs that represent concepts as high-dimensional embeddings. Our architecture, which we call Intervention-aware Concept Embedding Model (IntCEM), has two key distinctive features. First, it learns an intervention policy end-to-end which imposes a prior over which concepts an expert could be queried to intervene on next to maximise intervention receptiveness. Second, its loss function includes an explicit regulariser that penalises IntCEM for mispredicting the task label after a series of train-time interventions sampled from its learned policy. We highlight the following contributions:
@@ 2 Background and Previous Work
Methods in Concept-based Explainable AI Bau et al. (2017); Fong and Vedaldi (2018); Kim et al. (2018); Goyal et al. (2019a); Ghorbani et al. (2019); Kazhdan et al. (2020); Yeh et al. (2020) advocate for explaining DNNs by constructing explanations for their predictions using high-level concepts captured within their latent space. Within these methods, Concept Bottleneck Models (CBMs) Koh et al. (2020) provide a framework for developing concept-based interpretable DNNs.
@@
Because $f$ operates only on the concept scores predicted by $g$, if the set of training concepts is not predictive of the downstream task, then a CBM will be forced to choose between being highly accurate at predicting concepts or at predicting task labels. Concept Embedding Models (CEMs) Espinosa Zarlenga et al. (2022) are a generalisation of CBMs that address this trade-off using high-dimensional concept representations in their bottlenecks. This design allows information of concepts not provided during training to flow into the label predictor via two $m$-dimensional vector representations (i.e., embeddings) for each training concept $c_{i}$: $\mathbf{c}^{+}_{i}\in\mathbb{R}^{m}$, an embedding representing $c_{i}$ when it is active (i.e., $c_{i}$ = 1), and $\mathbf{c}^{-}_{i}\in\mathbb{R}^{m}$, an embedding representing $c_{i}$ when it is inactive (i.e., $c_{i}$ = 0).
When processing sample $\mathbf{x}$, for each training concept $c_{i}$ a CEM constructs $\mathbf{c}^{+}_{i}$ and $\mathbf{c}^{-}_{i}$ by feeding $\mathbf{x}$ into two learnable models $\phi_{i}^{+},\phi_{i}^{-}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ implemented as DNNs with a shared preprocessing module. These embeddings are then passed to a learnable scoring model $s:\mathbb{R}^{2m}\rightarrow[0,1]$, shared across all concepts, that predicts the probability $\hat{p}_{i}=s([\hat{\mathbf{c}}^{+}_{i},\hat{\mathbf{c}}^{-}_{i}])$ of concept $c_{i}$ being active. With these probabilities, one can define a CEM’s concept encoder $g(\mathbf{x})=\hat{\mathbf{c}}:=[\hat{\mathbf{c}}_{1},\cdots,\hat{\mathbf{c}}_{k}]\in\mathbb{R}^{km}$ by mixing the positive and negative embeddings of each concept to generate a final concept embedding $\hat{\mathbf{c}}_{i}:=\hat{p}_{i}\hat{\mathbf{c}}^{+}_{i}+(1-\hat{p}_{i})\hat{\mathbf{c}}^{-}_{i}$. Finally, a CEM can predict a label $\hat{\mathbf{y}}$ for sample $\mathbf{x}$ by passing $\hat{\mathbf{c}}$ (i.e., its “bottleneck”) to a learnable label predictor $f(\hat{\mathbf{c}})$ whose output can be explained with the concept probabilities $\hat{\mathbf{p}}:=[\hat{p}_{1},\cdots,\hat{p}_{k}]^{T}$. This generalisation of CBMs has been empirically shown to outperform CBMs especially when the set of concept annotations is incomplete Espinosa Zarlenga et al. (2022).
@@
We note that although we call the operation defined in Equation 1 a “concept intervention” to follow the term used in the concept learning literature, it should be distinguished from interventions in causal models Pearl et al. (2000); Goyal et al. (2019b); O’Shaughnessy et al. (2020). Instead, concept interventions in this context are connected, yet not identical, to previous work in active feature acquisition in which one has access to an expert at test-time to request the ground truth value of a small number of input features Shim et al. (2018); Li and Oliva (2021); Strauss and Oliva (2022).
@@
In practice, it may not be sensible to ask a human to intervene on every concept (Barker et al., 2023); instead, we may want a preferential ordering of which concepts to query first (e.g., some concepts may be highly informative of the output task, yet hard for the model to predict). “Intervention policies” have been developed to produce a sensible ordering for concept interventions Shin et al. (2023); Sheth et al. (2022); Chauhan et al. (2022). These orderings can be produced independently of a particular instance (i.e., a static policy), or selected on a per-instance basis (a dynamic policy). The latter takes as an input a mask of previously intervened concepts $\mathbf{\mu}$ and a set of predicted concept probabilities $\hat{\mathbf{p}}$, and determines which concept should be requested next from an expert, where the retrieved concept label is assumed to be “ground truth”. Two such policies are Cooperative Prediction (CooP) Chauhan et al. (2022) and Expected Change in Target Prediction (ECTP) Shin et al. (2023), which prioritise concept $c_{i}$ if intervening on $c_{i}$ next leads to the largest change in the probability of the currently predicted class in expectation (with the expectation taken over the distribution given by $p(c_{i}=1)=\hat{p}_{i}$). These policies, however, may become costly as the number of concepts increases and share the limitation that, by maximising the expected change in probability in the predicted class, they do not guarantee that the probability mass shifts towards the correct class.
@@ 3 Intervention-Aware Concept Embedding Models
Although there has been significant work on understanding how CBMs react to interventions Espinosa Zarlenga et al. (2022); Havasi et al. (2022); Chauhan et al. (2022); Sheth et al. (2022), the fact that these models are trained without any expert intervention – yet are expected to be highly receptive to expert feedback at test-time – is often overlooked. To the best of our knowledge, the only exception comes from Espinosa Zarlenga et al. (2022), wherein the authors introduced RandInt, a procedure that randomly intervenes on the model during train-time. However, RandInt assumes that a model should be equally penalised for mispredicting $y$ regardless of how many interventions have been performed. This leads to CEMs lacking the incentive to perform better when given more feedback. Moreover, RandInt assumes that all concepts are equally likely to be intervened on. This inadequately predisposes a model to assume that concepts will be randomly intervened on at test-time.
@@ 3.2 IntCEM’s Training Procedure
The purpose of $\mathcal{L}_{\text{roll}}$ is to learn an intervention policy that prioritises the selection of concepts for which the associated concept embeddings fails to represent the ground truth concept labels, i.e., select concepts which benefit from human intervention. We achieve this through a train-time “behavioural cloning” Bain and Sammut (1995); Ross et al. (2011) approach where $\psi$ is trained to learn to predict the action taken by an optimal greedy policy at each step. For this, we take advantage of the presence of ground truth concepts and task labels during training and incentivise $\psi$ to mimic “Skyline”, an oracle optimal policy proposed by Chauhan et al. (2022). Given a bottleneck $\hat{\mathbf{c}}$ and ground-truth labels $(y,\mathbf{c})$, the Skyline policy selects $c_{*}(\mathbf{x},\mathbf{\mu},\mathbf{c},y):=\operatorname*{arg\,max}_{\begin{subarray}{c}1\leq i\leq k\end{subarray}}f\big{(}\tilde{g}(\mathbf{x},\mathbf{\mu}\vee\mathds{1}_{i},\mathbf{c})\big{)}_{y}$ as its next concept, where $f(\cdot)_{y}$ is the probability of class $y$ predicted by $f$ and $\mathbf{\mu}\vee\mathds{1}_{i}$ represents the action of adding an intervention on concept $i$ to $\mathbf{\mu}$. Intuitively, $c_{*}$ corresponds to the concept whose ground-truth label and intervention would yield the highest probability over the ground-truth class $y$. We use such a demonstration to provide feedback to $\psi$ through the following loss:
@@
Here $\gamma\in[1,\infty)$ is a scaling factor that penalises IntCEM more heavily for mispredicting $y$ at the end of the trajectory, by a factor of $\gamma^{T}>1$, than for mispredicting $y$ at the start of the trajectory. In practice, we observe that a value in $\gamma\in[1.1,1.5]$ works well; we use $\gamma=1.1$ in our experiments unless specified otherwise. For an ablation study showing that IntCEM’s intervention performance surpasses that of existing methods for a wide array of scaling factors $\gamma$, see Appendix 6.6.
A key realisation is that expressing interventions as in Equation (1) yields an operator which is differentiable with respect to the concept intervention mask $\mathbf{\mu}$. This trick allows us to backpropagate gradients from $\mathcal{L}_{\text{pred}}$ into $\psi$ when the sampling operation $\mathbf{\eta}^{(t)}\sim p(\mathbf{\eta}|\psi(\hat{\mathbf{c}}^{(t-1)},\mathbf{\mu}^{(t-1)}))$ is differentiable with respect to its parameters. In this work, this is achieved by relaxing our trajectory sampling using a differentiable Gumbel-Softmax Jang et al. (2016) categorical sampler (see Appendix 6.2 for details).
@@
The outer expectation can be optimised via stochastic gradient descent, while the inner expectation can be estimated using Monte Carlo samples from the user-selected priors $(p(\mathbf{\mu}),p(T))$. For the sake of simplicity, we estimate the inner expectation using a single Monte Carlo sample per training step for both the initial mask $\mathbf{\mu}^{(0)}$ and the trajectory horizon $T$. In our experiments, we opt to use a Bernoulli prior for $p(\mathbf{\mu})$, where each concept is selected with probability $p_{\text{int}}=0.25$, and a discrete uniform prior $\text{Unif}(\{1,\cdots,T_{\text{max}}\})$ for $p(T)$, where $T_{\text{max}}$ is annealed within $[2,6]$. Although domain-specific knowledge can be incorporated into these priors, we leave this for future work and show in Appendix 6.6.3 that IntCEMs are receptive to interventions as we vary $p_{\text{int}}$ and $T_{\text{max}}$.
@@ 4 Experiments
We consider five vision tasks: (1) MNIST-Add, a task inspired by the UMNIST dataset Collins et al. (2023) where one is provided with 12 MNIST Deng (2012) images containing digits in $\{0,\cdots,9\}$, as well as their digit labels as concept annotations, and needs to predict if their sum is at least half of the maximum attainable, (2) MNIST-Add-Incomp, a concept-incomplete Yeh et al. (2020) extension of the MNIST-Add task where only 8/12 operands are provided as concept annotations, (3) CUB Wah et al. (2011), a bird classification task whose set of concepts cover 112 bird features, e.g., “wing shape”, selected by Koh et al. (2020), (4) CUB-Incomp, a concept-incomplete extension of CUB where we provide only $25\%$ of all concept groups, and (5) CelebA, a task on the Celebrity Attributes Dataset Liu et al. (2018) whose task and concept labels follow those selected in Espinosa Zarlenga et al. (2022). For more details refer to Appendix 6.3.
@@
We evaluate IntCEMs against CEMs with the exact same architecture, averaging all metrics of interest over five different random initialisations. We also include CBMs that are trained jointly (Joint CBM), sequentially (Seq. CBM) and independently (Ind. CBM) as part of our evaluation to be able to compare IntCEM’s policy against state-of-art intervention policies developed for CBMs. Since previous work shows that interventions in joint CBMs are sensitive to its bottleneck activation function Koh et al. (2020); Espinosa Zarlenga et al. (2023), we include joint CBMs with a sigmoidal (Joint CBM-Sigmoid) and logit-based (Joint CBM-Logit) bottleneck. We note that we do not include other noteworthy concept-based models such as leakage-free CBMs Havasi et al. (2022), GlanceNets Marconato et al. (2022), Self-explaining Neural Networks (SENNs) Alvarez-Melis and Jaakkola (2018), Concept Whitening (CW) Chen et al. (2020), or Concept Model Extraction (CME) Kazhdan et al. (2020) as these either lack a well-defined mechanism to be intervened on (e.g., as in SENNs and CW) or represent more constrained versions of other baselines such as CEMs (e.g., leakage-free CBMs, GlanceNets, and CME).
Throughout our experiments, we strive for a fair comparison by using the same architectures for the concept encoder $g$ and label predictor $f$ of all baselines. Unless specified otherwise, for each task, we tune IntCEM’s $\lambda_{\text{roll}}$ hyperparameter by varying it over $\{5,1,0.1\}$ and selecting the model with the highest area under the curve defined by plotting the validation task accuracy vs number of interventions. For the CUB-based tasks and CelebA, we use the same values of $\lambda_{concept}$ selected in Espinosa Zarlenga et al. (2022). Otherwise, we select this value from $\{10,5,1\}$ using CEM’s validation error. Following Espinosa Zarlenga et al. (2022), we use embeddings with $m=16$ activations and a value of $p_{\text{int}}=0.25$ for CEM’s RandInt. For further details on model and training hyperparameters, see Appendix 6.4.

### v3

@@ 1 Introduction
It is important to know how to ask for help, but also important to know how to receive help. Knowing how to react to feedback positively allows for bias correction [1] and efficient learning [2] while being instrumental for mass collaborations [3]. Nevertheless, although the uptake of feedback is ubiquitous in real-world decision-making, the same cannot be said about modern artificial intelligence (AI) systems, where deployment tends to be in isolation from experts whose feedback could be queried.
Progress in this aspect has recently come from Explainable AI, where interpretable Deep Neural Networks (DNNs) that can benefit from expert feedback at test-time have been proposed [4, 5, 6, 7]. In particular, Concept Bottleneck Models (CBMs) [4], a family of interpretable DNNs, enable expert-model interactions by generating, as part of their inference process, explanations for their predictions using high-level “concepts”. Such explanations allow human experts to better understand a CBM’s prediction based on interpretable units of information (e.g., input is a “cat” because it has “paws” and “whiskers”) rather than low-level features (e.g., input pixels). This design enables humans to provide feedback to the CBM at test-time via concept interventions [4] (Figure 1), a process specific to CBM-like architectures in which an expert analyses a CBM’s inferred concepts and corrects mispredictions. Concept interventions allow a CBM to update its final prediction after incorporating new expert feedback, and have been shown to lead to significant performance boosts [4, 5, 8, 9]. Nevertheless, and in contrast to these results, recent works have found that concept interventions can potentially increase a CBM’s test-time error depending on its choice of architecture [10] and its set of train-time annotations [5]. This discrepancy suggests that a CBM’s receptiveness to interventions is neither guaranteed nor fully understood.
@@
In this paper, we posit that the sensitivity of a CBM’s receptiveness to interventions is an artefact of its training objective function. Specifically, we argue that CBMs lack an explicit incentive to be receptive to interventions as they are neither exposed to interventions at train-time nor optimised to perform well under interventions. To address this, we propose a novel intervention-aware objective function and architecture based on Concept Embedding Models (CEMs) [5], a generalisation of CBMs that represent concepts as high-dimensional embeddings. Our architecture, which we call Intervention-aware Concept Embedding Model (IntCEM), has two key distinctive features. First, it learns an intervention policy end-to-end which imposes a prior over which concepts an expert could be queried to intervene on next to maximise intervention receptiveness. Second, its loss function includes an explicit regulariser that penalises IntCEM for mispredicting the task label after a series of train-time interventions sampled from its learned policy. We highlight the following contributions:
@@ 2 Background and Previous Work
Methods in Concept-based Explainable AI [11, 12, 13, 14, 15, 16, 17] advocate for explaining DNNs by constructing explanations for their predictions using high-level concepts captured within their latent space. Within these methods, Concept Bottleneck Models (CBMs) [4] provide a framework for developing concept-based interpretable DNNs.
@@
Because $f$ operates only on the concept scores predicted by $g$, if the set of training concepts is not predictive of the downstream task, then a CBM will be forced to choose between being highly accurate at predicting concepts or at predicting task labels. Concept Embedding Models (CEMs) [5] are a generalisation of CBMs that address this trade-off using high-dimensional concept representations in their bottlenecks. This design allows information of concepts not provided during training to flow into the label predictor via two $m$-dimensional vector representations (i.e., embeddings) for each training concept $c_{i}$: $\mathbf{c}^{+}_{i}\in\mathbb{R}^{m}$, an embedding representing $c_{i}$ when it is active (i.e., $c_{i}$ = 1), and $\mathbf{c}^{-}_{i}\in\mathbb{R}^{m}$, an embedding representing $c_{i}$ when it is inactive (i.e., $c_{i}$ = 0).
When processing sample $\mathbf{x}$, for each training concept $c_{i}$ a CEM constructs $\mathbf{c}^{+}_{i}$ and $\mathbf{c}^{-}_{i}$ by feeding $\mathbf{x}$ into two learnable models $\phi_{i}^{+},\phi_{i}^{-}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ implemented as DNNs with a shared preprocessing module. These embeddings are then passed to a learnable scoring model $s:\mathbb{R}^{2m}\rightarrow[0,1]$, shared across all concepts, that predicts the probability $\hat{p}_{i}=s([\hat{\mathbf{c}}^{+}_{i},\hat{\mathbf{c}}^{-}_{i}])$ of concept $c_{i}$ being active. With these probabilities, one can define a CEM’s concept encoder $g(\mathbf{x})=\hat{\mathbf{c}}:=[\hat{\mathbf{c}}_{1},\cdots,\hat{\mathbf{c}}_{k}]\in\mathbb{R}^{km}$ by mixing the positive and negative embeddings of each concept to generate a final concept embedding $\hat{\mathbf{c}}_{i}:=\hat{p}_{i}\hat{\mathbf{c}}^{+}_{i}+(1-\hat{p}_{i})\hat{\mathbf{c}}^{-}_{i}$. Finally, a CEM can predict a label $\hat{\mathbf{y}}$ for sample $\mathbf{x}$ by passing $\hat{\mathbf{c}}$ (i.e., its “bottleneck”) to a learnable label predictor $f(\hat{\mathbf{c}})$ whose output can be explained with the concept probabilities $\hat{\mathbf{p}}:=[\hat{p}_{1},\cdots,\hat{p}_{k}]^{T}$. This generalisation of CBMs has been empirically shown to outperform CBMs especially when the set of concept annotations is incomplete [5].
@@
We note that although we call the operation defined in Equation 1 a “concept intervention” to follow the term used in the concept learning literature, it should be distinguished from interventions in causal models [18, 19, 20]. Instead, concept interventions in this context are connected, yet not identical, to previous work in active feature acquisition in which one has access to an expert at test-time to request the ground truth value of a small number of input features [21, 22, 23].
@@
In practice, it may not be sensible to ask a human to intervene on every concept [24]; instead, we may want a preferential ordering of which concepts to query first (e.g., some concepts may be highly informative of the output task, yet hard for the model to predict). “Intervention policies” have been developed to produce a sensible ordering for concept interventions [8, 25, 9]. These orderings can be produced independently of a particular instance (i.e., a static policy), or selected on a per-instance basis (a dynamic policy). The latter takes as an input a mask of previously intervened concepts $\mathbf{\mu}$ and a set of predicted concept probabilities $\hat{\mathbf{p}}$, and determines which concept should be requested next from an expert, where the retrieved concept label is assumed to be “ground truth”. Two such policies are Cooperative Prediction (CooP) [9] and Expected Change in Target Prediction (ECTP) [8], which prioritise concept $c_{i}$ if intervening on $c_{i}$ next leads to the largest change in the probability of the currently predicted class in expectation (with the expectation taken over the distribution given by $p(c_{i}=1)=\hat{p}_{i}$). These policies, however, may become costly as the number of concepts increases and share the limitation that, by maximising the expected change in probability in the predicted class, they do not guarantee that the probability mass shifts towards the correct class.
@@ 3 Intervention-Aware Concept Embedding Models
Although there has been significant work on understanding how CBMs react to interventions [5, 6, 9, 25], the fact that these models are trained without any expert intervention – yet are expected to be highly receptive to expert feedback at test-time – is often overlooked. To the best of our knowledge, the only exception comes from Espinosa Zarlenga et al. [5], wherein the authors introduced RandInt, a procedure that randomly intervenes on the model during train-time. However, RandInt assumes that a model should be equally penalised for mispredicting $y$ regardless of how many interventions have been performed. This leads to CEMs lacking the incentive to perform better when given more feedback. Moreover, RandInt assumes that all concepts are equally likely to be intervened on. This inadequately predisposes a model to assume that concepts will be randomly intervened on at test-time.
@@ 3.2 IntCEM’s Training Procedure
The purpose of $\mathcal{L}_{\text{roll}}$ is to learn an intervention policy that prioritises the selection of concepts for which the associated concept embeddings fails to represent the ground truth concept labels, i.e., select concepts which benefit from human intervention. We achieve this through a train-time “behavioural cloning” [26, 27] approach where $\psi$ is trained to learn to predict the action taken by an optimal greedy policy at each step. For this, we take advantage of the presence of ground truth concepts and task labels during training and incentivise $\psi$ to mimic “Skyline”, an oracle optimal policy proposed by Chauhan et al. [9]. Given a bottleneck $\hat{\mathbf{c}}$ and ground-truth labels $(y,\mathbf{c})$, the Skyline policy selects $c_{*}(\mathbf{x},\mathbf{\mu},\mathbf{c},y):=\operatorname*{arg\,max}_{\begin{subarray}{c}1\leq i\leq k\end{subarray}}f\big{(}\tilde{g}(\mathbf{x},\mathbf{\mu}\vee\mathds{1}_{i},\mathbf{c})\big{)}_{y}$ as its next concept, where $f(\cdot)_{y}$ is the probability of class $y$ predicted by $f$ and $\mathbf{\mu}\vee\mathds{1}_{i}$ represents the action of adding an intervention on concept $i$ to $\mathbf{\mu}$. Intuitively, $c_{*}$ corresponds to the concept whose ground-truth label and intervention would yield the highest probability over the ground-truth class $y$. We use such a demonstration to provide feedback to $\psi$ through the following loss:
@@
Here $\gamma\in[1,\infty)$ is a scaling factor that penalises IntCEM more heavily for mispredicting $y$ at the end of the trajectory, by a factor of $\gamma^{T}>1$, than for mispredicting $y$ at the start of the trajectory. In practice, we observe that a value in $\gamma\in[1.1,1.5]$ works well; we use $\gamma=1.1$ in our experiments unless specified otherwise. For an ablation study showing that IntCEM’s intervention performance surpasses that of existing methods for a wide array of scaling factors $\gamma$, see Appendix A.6.
A key realisation is that expressing interventions as in Equation (1) yields an operator which is differentiable with respect to the concept intervention mask $\mathbf{\mu}$. This trick allows us to backpropagate gradients from $\mathcal{L}_{\text{pred}}$ into $\psi$ when the sampling operation $\mathbf{\eta}^{(t)}\sim p(\mathbf{\eta}|\psi(\hat{\mathbf{c}}^{(t-1)},\mathbf{\mu}^{(t-1)}))$ is differentiable with respect to its parameters. In this work, this is achieved by relaxing our trajectory sampling using a differentiable Gumbel-Softmax [28] categorical sampler (see Appendix A.2 for details).
@@
The outer expectation can be optimised via stochastic gradient descent, while the inner expectation can be estimated using Monte Carlo samples from the user-selected priors $(p(\mathbf{\mu}),p(T))$. For the sake of simplicity, we estimate the inner expectation using a single Monte Carlo sample per training step for both the initial mask $\mathbf{\mu}^{(0)}$ and the trajectory horizon $T$. In our experiments, we opt to use a Bernoulli prior for $p(\mathbf{\mu})$, where each concept is selected with probability $p_{\text{int}}=0.25$, and a discrete uniform prior $\text{Unif}(\{1,\cdots,T_{\text{max}}\})$ for $p(T)$, where $T_{\text{max}}$ is annealed within $[2,6]$. Although domain-specific knowledge can be incorporated into these priors, we leave this for future work and show in Appendix A.6.3 that IntCEMs are receptive to interventions as we vary $p_{\text{int}}$ and $T_{\text{max}}$.
@@ 4 Experiments
We consider five vision tasks: (1) MNIST-Add, a task inspired by the UMNIST dataset [29] where one is provided with 12 MNIST [30] images containing digits in $\{0,\cdots,9\}$, as well as their digit labels as concept annotations, and needs to predict if their sum is at least half of the maximum attainable, (2) MNIST-Add-Incomp, a concept-incomplete [17] extension of the MNIST-Add task where only 8/12 operands are provided as concept annotations, (3) CUB [31], a bird classification task whose set of concepts cover 112 bird features, e.g., “wing shape”, selected by Koh et al. [4], (4) CUB-Incomp, a concept-incomplete extension of CUB where we provide only $25\%$ of all concept groups, and (5) CelebA, a task on the Celebrity Attributes Dataset [32] whose task and concept labels follow those selected in [5]. For more details refer to Appendix A.3.
@@
We evaluate IntCEMs against CEMs with the exact same architecture, averaging all metrics of interest over five different random initialisations. We also include CBMs that are trained jointly (Joint CBM), sequentially (Seq. CBM) and independently (Ind. CBM) as part of our evaluation to be able to compare IntCEM’s policy against state-of-art intervention policies developed for CBMs. Since previous work shows that interventions in joint CBMs are sensitive to its bottleneck activation function [4, 10], we include joint CBMs with a sigmoidal (Joint CBM-Sigmoid) and logit-based (Joint CBM-Logit) bottleneck. We note that we do not include other noteworthy concept-based models such as leakage-free CBMs [6], GlanceNets [33], Self-explaining Neural Networks (SENNs) [34], Concept Whitening (CW) [35], or Concept Model Extraction (CME) [16] as these either lack a well-defined mechanism to be intervened on (e.g., as in SENNs and CW) or represent more constrained versions of other baselines such as CEMs (e.g., leakage-free CBMs, GlanceNets, and CME).
Throughout our experiments, we strive for a fair comparison by using the same architectures for the concept encoder $g$ and label predictor $f$ of all baselines. Unless specified otherwise, for each task, we tune IntCEM’s $\lambda_{\text{roll}}$ hyperparameter by varying it over $\{5,1,0.1\}$ and selecting the model with the highest area under the curve defined by plotting the validation task accuracy vs number of interventions. For the CUB-based tasks and CelebA, we use the same values of $\lambda_{concept}$ selected in [5]. Otherwise, we select this value from $\{10,5,1\}$ using CEM’s validation error. Following [5], we use embeddings with $m=16$ activations and a value of $p_{\text{int}}=0.25$ for CEM’s RandInt. For further details on model and training hyperparameters, see Appendix A.4.